BRICS Notes Series, Abstracts, 2005

December 23, 2005

This document is also available as PostScript and DVI.

References

[NS-05-6] PDF.

Anders Møller, editor. PLAN-X 2006 Informal Proceedings, (Charleston, South Carolina,

January 14, 2006), December 2005. ii+92. Abstract: The workshop aims at providing a meeting ground for

researchers from the XML, programming language, and database communities. XML is already a de-facto industry standard for data exchange, it has from an early stage been embraced by database researchers, and it is gaining increasing interest from programming language researchers. At this workshop we hope to present recent results, identify new challenges, and inspire the programming language community to focus on XML. The focus of the workshop is on methods, tools, and theories for processing XML. Example topics include (but are not limited to) XML parsing, XML type system and schemas, analysis and implementation of technologies such as XPath, XSLT, and XQuery, and integration of XML in both general-purpose and domain-specific programming languages.

[NS-05-5] PDF.

Patrick Cousot, Lisbeth Fajstrup, Eric Goubault, Maurice Herlihy,

Kim G. Larsen, and Martin Raußen, editors. Preliminary Proceedings of the Workshop on Geometry and Topology

in Concurrency, GETCO ’05, (San Francisco, California, USA, August 21, 2005), August 2005.

vi+44. Abstract: This volume contains the preliminary proceedings of

GETCO 2005 held as a satellite workshop of CONCUR 2005.

[NS-05-4] PDF.

Scott A. Smolka and Jiří Srba, editors. Preliminary Proceedings of the 7th International Workshop on Veri-

fication of Infinite-State Systems, INFINITY ’05, (San Francisco, USA, August 27, 2005), June 2005. vi+64 pp. Abstract: This volume contains the preliminary proceedings of IN-

FINITY’05 held in San Francisco, USA, August 27, 2005. The final proceedings will appear as a volume in the ENTCS series.

[NS-05-3] PDF.

Luca Aceto and Andrew D. Gordon, editors. Short Contributions from the Workshop on Algebraic Process Calculi:

The First Twenty Five Years and Beyond, PA ’05, (Bertinoro, Forlì, Italy, August 1–5, 2005), June 2005.

vi+239 pp. Abstract: This volume contains short contributions from the work-

shop on “Algebraic Process Calculi: The First Twenty Five Years and Beyond”, held in the period August 1–5, 2005, at the University Residential Centre of Bertinoro, Forlì, Italy.

[NS-05-2] PostScript, PDF, DVI.

Luca Aceto and Willem Jan Fokkink. The Quest for Equational Axiomatizations of Parallel Composition:

Status and Open Problems. May 2005. 7 pp. To appear in a volume of the BRICS Notes Series devoted

to the workshop “Algebraic Process Calculi: The First Twenty Five Years and Beyond”, August 1–5, 2005, University of Bologna Residential Center Bertinoro (Forlì), Italy.

Abstract: This essay recounts the story of the quest for equational axiomatizations of parallel composition operators in process description languages, and of similar results in the classic field of formal language theory. Some of the outstanding open problems are also mentioned.

[NS-05-1] Luca Aceto, Magnus Mar Halldorsson, and Anna Ingólfsdóttir. What is Theoretical Computer Science? April 2005. 13 pp. Abstract: This article presents a bird’s eye view of Theoretical Com-

puter Science aimed at a general scientific audience. It then focuses on a selected area within this general field, and discusses some of the applications of results from that area and of the fundamental theoretical questions that drive its development.